OpenStudy (anonymous):

Can someone please help me and show me how to solve this problem step by step? Find a circle with the center at (7,1), graph contains the point (1,-7).

OpenStudy (anonymous):

The standard form equation of a circle is a way to express the definition of a circle on the coordinate plane. On the coordinate plane, the formula becomes (x−h)^2+(y−k)^2=r^2 h and k are the x and y coordinates of the center of the circle (x−9)^2+(y−6)^2=100 is a circle centered at (9,6) with a radius of 10

OpenStudy (anonymous):

Now you know where to center your circle, so you need to figure out your radius to include the point you need. You got it?

OpenStudy (anonymous):

Yeah I understand what you've said so far, I'm just not sure how to find the radius.

OpenStudy (anonymous):

The equation of this problem would be (x-7)^2+(y-1)^2=r^2

OpenStudy (anonymous):

Right, now try drawing a picture. You need the distance from the center to the point asked for, thats your radius.

OpenStudy (anonymous):

So I draw a circle on a graph. With the center at (7,1) and plot the point given, (1,-7). Correct?

OpenStudy (anonymous):